2. A normal dose of ibuprofen is 200. mg. How many ibuprofen molecules are there in a dose of ibuprofen? The molecular formula of ibuprofen is C13H18O2. (Hint: grams to mol and mol to molecules)


3-Propane fuel is used in BBQs. Use the following reaction to determine how many moles of carbon dioxide gas are produced from burning of 6 moles of propane gas.
C3H8(g) + 5O2(g) ®  4H2O(g) + 3CO2(g)


4- A. If a person inhales 51 moles of oxygen, how many grams of oxygen molecules are being inhaled?

B. How many molecules are being inhaled?
Answer #1

2.
The molecular formula of ibuprofen is C13H18O2
The atomic masses of C, H and O are 12 g/mol, 1 g/mol and 16 g/mol respectively.
The molar mass of ibuprofen is 13 (12) + 18(1) + 2 (16) = 206 \: g/mol
200 mg = 0.200 g
Number of moles = \dfrac {0.200 \: g}{206 \: g/mol} = 0.0009709 \: mol
Number of molecules = 0.0009709 \: mol \times 6.023 \times 10^{23} = 5.85 \times 10^{20} \; molecules

3.
The burning of 1 mole propane produces 3 moles of CO2.
The burning of 6 mole propane produces 3 \times 6 = 18 moles of CO2

4.A
The molar mass of oxygen is 32 g/mol.
Grams of oxygen molecules inhaled = 51 \: mol \times 32 \: g/mol = 1632 \: g
B
Number of molecules being inhaled = 51 \: moles \times 6.023 \times 10^{23} \: molecules/mol = 3.07 \times 10^{25} \: molecules